Find a polynomial of degree 4 with -2 as a zero of multiplicity 2 and 0 and 3 as zeros of multiplicity 1.
Answer by lwsshak3(6505)

Find a polynomial of degree 4 with -2 as a zero of multiplicity 2 and 0 and 3 as zeros of multiplicity 1.
..
Factors:
(x+2)(x+2)(x)(x-3)
=(x^2+4x+4)(x)(x-3)
=(x^3+4x^2+4x)(x-3)
=x^4+4x^3+4x^2-3x^3-12x^2-12x
=x^4+x^3-8x^2-12x (ans)